OpenStudy (anonymous):

simplify sqrt{300x^16}

OpenStudy (anonymous):

we can use the rule of sqrt(a*b) = sqrt(a)*sqrt(b) so in this case we will get sqrt(300)*sqrt(x^16) = sqrt(300)*x^4 and we know 300=100*3 so we now will get sqrt(100)*sqrt(3)*x^4 = 10x^4*sqrt(3) :)

OpenStudy (unklerhaukus):

\[\sqrt{300x^{16}}=\sqrt{3\times10^2\times(x^8)^2}\]

jhonyy9 (jhonyy9):

hamza check it newly because not is right

jhonyy9 (jhonyy9):

hamza_b23 chek it your answer please because not is right

OpenStudy (unklerhaukus):

\[\sqrt{300x^{16}}=\sqrt{3\times10^2\times(x^8)^2}=10x^8\sqrt{3}\]

OpenStudy (unklerhaukus):

\[(x^n)^m=x^{nm}\]